# Hofstadter-Toda spectral duality and quantum groups

### v1

@@
Abstract.
The Hofstadter model allows to describe and understand several phenomena in condensed matter such as the quantum Hall effect, Anderson localization, charge pumping, and flat-bands in quasiperiodic structures, and is a rare example of fractality in the quantum world. An apparently unrelated system, the relativistic Toda lattice, has been extensively studied in the context of complex nonlinear dynamics, and more recently for its connection to supersymmetric Yang–Mills theories and topological string theories on Calabi–Yau manifolds in high-energy physics. Here we discuss a recently discovered spectral relationship between the Hofstadter model and the relativistic Toda lattice which has been later conjectured to be related to the Langlands duality of quantum groups. Moreover, by employing similarity transformations compatible with the quantum group structure, we establish a formula parametrizing the energy spectrum of the Hofstadter model in terms of elementary symmetric polynomials and Chebyshev polynomials. The main tools used are the spectral duality of tridiagonal matrices and the representation theory of the elementary quantum group.
Key words and phrases:
quantum Hall effect, Hofstadter model, Aubry–André model, quantum groups, Toda lattice.
2020 Mathematics Subject Classification:
81R50; 82D03, 81R60
@@ Introduction
Among the diversity of natural phenomena, intriguing connections sometimes emerge, where different physical systems share similar mathematical equations or structures. These unexpected parallels may seem coincidental or, conversely, hint at a deeper unity within the physical and mathematical worlds. One such example is the relationship between the relativistic Toda lattice model [Tod67, Rui90] in high energy physics and the Hofstadter [Har55, Hof76, AA80] model in condensed matter physics.
@@
The Toda lattice model describes a one-dimensional lattice of $N$ particles with exponentially decaying interactions [Tod67]. Initially conceived as a simple toy model for one-dimensional crystals, it is a prominent example of a nonlinear system in mathematical physics, which stands out for the remarkable property of being exactly solvable and integrable. Moreover, the model exhibits the emergence of solitonic excitations, providing insights into the complex dynamics of nonlinear systems. Furthermore, its relativistic generalization [Rui90] is studied in high-energy physics for its connection to supersymmetric Yang–Mills theories in five dimensions [Nek98] and topological string theories on Calabi–Yau manifolds [KKV97, ${}^{+}$12, HM16, HKT16, HSX17].
@@
On the other hand, the Hofstadter model describes noninteracting charged fermions (e.g., electrons) on a two-dimensional lattice in a magnetic field (more generally, a gauge field) perpendicular to the plane or, alternatively, noninteracting fermions in a one-dimensional quasiperiodic lattice [Har55, Hof76, AA80]. The most distinctive feature of the model is the presence of a fractal energy spectrum, which is one of the few examples of fractals in quantum physics, and the very rich phase diagram [OA01] with topological gapped phases indexed by a topological invariant, the Chern number of the occupied bands [TKNdN82]. The Hofstadter model is the discrete counterpart of a quantum Hall system, in the sense that it describes the Landau levels of free electrons in a magnetic field regularized on a discrete lattice [TKNdN82]. Indeed, in the weak field limit, i.e., when the magnetic cyclotron radius becomes much larger than the lattice constant, the quantum Hall system and the Hofstadter model become physically equivalent. It is however when the magnetic cyclotron radius is comparable with the lattice constant that the fractal properties emerge, as a consequence of the incommensuration between these two characteristic lengths.
@@
The Hofstadter Hamiltonian is moreover identical to the model introduced by Aubry and André [AA80], describing fermions on a one-dimensional discrete lattice in the presence of a harmonic potential not necessarily commensurate with the lattice parameter. In the Aubry–André model, the role of the magnetic flux per unit cell in the two-dimensional system is played by the angular wavenumber of the harmonic potential (in units of the lattice parameter), and the role of the second dimension is played by the phase shift of the harmonic potential with respect to the discrete lattice, giving rise to a so-called second synthetic dimension. This also leads to the emergence of topological charge pumping [Tho83], which shares the same topological origin as the quantum Hall effect. The fractal nature of the spectra emerges here as a consequence of the incommensuration between the periodicity of the harmonic potential and the periodicity of the underlying lattice. This model can describe the effects of disorder, including Anderson localization [AA80] and quasiperiodicity [${}^{+}$12, KRZ13, MN20].
@@
Physical realizations of the Hofstadter model include microwave photonic systems [KS98], Moiré superlattices in graphene [${}^{+}$13, ${}^{+}$13, ${}^{+}$13], ultracold atoms in incommensurate optical lattices [${}^{+}$13], and interacting phonons in superconducting qubits [${}^{+}$17].
@@
Although originating from distinct branches of physics, the $N=2$ relativistic Toda lattice and Hofstadter models are described by Hamiltonians which transform one into the other by a formal substitution of the position and momenta operators $x,p\mapsto\mathrm{i}x,\mathrm{i}p$. Mathematically, this corresponds to substituting Mathieu ($\cos$) and the modified Mathieu ($\cosh$) potentials or operators. Recently, a deep relationship between the energy spectra of the Hofstadter model, the relativistic Toda lattice model, and their modular dual was found by Hatsuda, Katsura, and Tachikawa [HKT16]. Building on these results, it has been conjectured that the Hofstadter model is intimately related to the Langlands duality of quantum groups [Ike18], suggesting a deep connection between the Langlands program, quantum geometry, and quasiperiodicity.
@@
We notice that both the Hofstadter and Toda lattice models are known to be related to the root systems $A_{N-1}$. Hence it should be no surprise that $\mathcal{U}_{q}(\mathfrak{sl}_{2})$ (the simplest quantum group) plays a central role in the study of these models. More generally, we stress the fact that noncommutative geometry is the most natural language to investigate the properties of the Hofstadter model and the related Aubry and André model, as well as other quasiperiodic and aperiodic structures emerging in condensed matter physics, such as quasicrystals, Penrose tilings, and disordered systems [Bel03, Put10, LH11, OOP11].
It should be emphasized that the first result below (Theorem 2.2) has already been understood [HKT16] in the context of the relativistic Toda model, and our contribution is a clarification of the relationship with the characteristic polynomials associated with the Hofstadter model, making use of a “quantum group-adjusted” gauge introduced by Wiegmann and Zabrodin [WZ94] and the spectral duality developed by Molinari (see [Mol97], where general Hamiltonian matrices similar to the one considered here are investigated).
The formula given in Theorem 3.9 and the parametrization in Theorem 4.2 are new, and they are obtained by using the representation theory of $\mathcal{U}_{q}(\mathfrak{sl}_{2})$ and standard formulas involving Chebyshev polynomials.
@@
The polynomial relationship of Eq. (1) was first proved by Hatsuda, Katsura, and Tachikawa [HKT16] by comparing the eigenvalue equation of the Hofstadter Hamiltonian $H$, the Toda Hamiltonian $H_{\text{Toda}}$, and its modular dual.
Here, we clarify the argument by using the spectral duality of tridiagonal matrices with corners developed in [Mol97] and proving that the polynomials involved are indeed the characteristic polynomials determining the spectrum of the Hofstadter model (as observed in [HKT16, p. 10]).
@@
For more details on the result above, see Theorem 4.2 and related discussion above it. This result should be interpreted as follows: in order to gain more insight into the mapping $E\mapsto\tilde{E}$, we can try to obtain explicit formulas for $f$ for all values of $\alpha=P/Q$ and then invert one side of Eq. (1) (at least locally) to get an analytic expression $\tilde{E}=\tilde{E}(E)$. This theorem is an attempt towards this goal. It should be noted that a formula for the Hofstadter spectrum at the mid-band point appeared previously in [Kre93], by using a strategy related to the one employed in this paper, but without involving the notion of quantum group.
@@
1. Preliminaries on mathematical structures and models
@@
There are several such constructions, most notably the Gell-Mann matrices and Sylvester’s shift and clock matrices [Syl11]. The former are Hermitian and traceless (just like the Pauli matrices), while the latter are unitary and traceless, which makes them preferable for our goals.
Definition 1.1.
@@
which justifies the name Weyl–Heisenberg matrices for $V$ and $U$. The Hofstadter model is described by the Hamiltonian $H=V+V^{*}+U+U^{*}$. Before going further, however, we need to introduce some additional mathematical structures. The relation in Eq. (2) defines a universal $C^{*}$-algebra $A_{\omega}$ studied by Rieffel [Rie81], known as noncommutative torus or rotation algebra. It is arguable that the naming “noncommutative torus” should be reserved for the case $\omega=\exp(2\pi\mathrm{i}\alpha)$ and $\alpha$ is irrational, unlike what we assumed so far. This is because when $\alpha$ is rational, $A_{\omega}$ is Morita equivalent to the commutative algebra of continuous functions on the standard torus. By contrast, when $\alpha$ is irrational, the Morita equivalence class is determined by the orbit of the modular action of $\mathrm{PGL}_{2}(\mathbb{Z})$ by linear fractional transformations on $\alpha$. Being these generated by
@@
in the mathematical physics literature $A_{\omega}$ is said to be in modular duality with $A_{\tilde{\omega}}$, where $\tilde{\omega}=\exp(2\pi\mathrm{i}\alpha^{-1})$. Let us remark here that the operator algebras community does not use this naming convention. However, other related dualities, such as Spanier–Whitehead duality in $K$-theory [Con94, NP20] or the property of being each other’s von Neumann algebra commutant (with coupling constant $\alpha$) [Fad95], are more frequently mentioned for $A_{\omega}$ and $A_{\tilde{\omega}}$.
Equation 3 provides our second contact point with the Hofstadter model, as those two transformations also appear as the fundamental self-similarity relations of the Hofstadter butterfly [Hof76, HKT16, Ike18]. In addition, the modular duality discussed above has inspired Faddeev’s notion of modular double of a quantum group [Fad14]. The modular double of $\mathcal{U}_{q}(\mathfrak{sl}_{2})$ plays a central role in the study of the relativistic Toda lattice [KLSTS02], particularly in the argument of [HKT16] where the polynomial relation which inspired this paper is established. To understand this, let us first recall the definition of the elementary quantum group.
Definition 1.2.
@@
Following [KLSTS02], we operate a Wick rotation on $\rho_{0},\tilde{\rho}_{0}$, by changing $T_{s}$ into $T_{\mathrm{i}s}$ and $S_{-\mathrm{i}s}$ into $S_{s}$. Note there is a fundamental difference between the representations of the quantum group and those of the rotation algebra: the former are obtained from unbounded representations of the latter. The newly obtained representations will be denoted $\rho,\tilde{\rho}$. Since they still commute, we can combine them with $\phi$ to form a representation of the modular double quantum group
@@
The general definition of this object involves the Langlands dual Lie algebra [KLSTS02]. However, we do not need to worry about such details here as $\mathfrak{sl}_{2}$ is self-dual. We can anticipate that the Toda Hamiltonian is $H\propto\rho\phi(E+F)$.
In condensed matter physics, the $N$-particle Toda lattice model [Tod67] is a seminal toy model describing massive particles on a one-dimensional lattice subject to a nearest-neighbor and exponentially decaying two-body interaction. In the periodic case (particles on a circle) the Hamiltonian reads
@@
with the periodic boundary condition $q_{0}\equiv q_{N}$, $q_{N+1}\equiv q_{1}$, where $p_{n}$ are the canonical momenta and $q_{n}$ the canonical coordinates satisfying the canonical Poisson bracket relations $\{q_{n},p_{m}\}=\delta_{nm}$. This Hamiltonian is nonrelativistic, being invariant up to Galilean transformations but not transformations of the Poincaré group, and is an early example of the family of integrable one-dimensional many-body problems, also called Calogero-Moser systems. A general approach to obtain a relativistic generalization of the Toda lattice model (as well as other Calogero-Moser systems) is to deform the nonrelativistic Hamiltonian to make it compatible with the Poincaré group (see [Rui90] for more details). This involves maintaining the nearest-neighbor exponential form of the interaction term, exponentiating the canonical momenta, and combining the resulting terms into an Hamiltonian which will coincide with the time-translation generator of the Poincaré group. Moreover, we want our Hamiltonian to be quantum, that is, to elevate the classical variables to quantum operators satisfying the canonical commutation relation $[q_{n},p_{m}]=\mathrm{i}\hbar\delta_{nm}$. Following this general recipe and using the same approach as [KLSTS02, HM16], we define our $N$-particle relativistic Toda lattice model as
@@
and also gives $[x,p]=\mathrm{i}\hbar/(2\pi\alpha)$. The relationship between quantum mechanics and quantum geometry is thus revealed if one sets $\hbar=2\pi\alpha$, which leads us back to $[x,p]=\mathrm{i}$ and $q=\mathrm{e}^{\mathrm{i}\pi\alpha}$. This Hamiltonian hence is expressed in terms of $\rho(v)$ and $\rho(vu)$, namely the first two generators of the generalized Clifford algebra introduced above (with the difference that these are infinite-dimensional operators). Finally, to uniform our notations to that of [HKT16], we introduce the canonical transformation $2\pi\alpha x^{\prime}=2\pi\alpha x+p$ (preserving the commutation relation $[x^{\prime},p]=[x,p]=\mathrm{i}$). By doing so and neglecting the prime, this yields the usual form of the relativistic Toda Hamiltonian for $N=2$ particles:
@@
The Hofstadter Hamiltonian describes the Hamiltonian of a charged particle in a magnetic field and in a periodic potential $V(x,y)$ describing a two-dimensional lattice. There are two opposite ways to derive this Hamiltonian: In the limit of strong fields, one takes the Hamiltonian of a charged particle in a field (describing the quantum Hall effect), which is diagonalized in terms of Landau levels, and then treats the periodic potential $V(x,y)$ as a perturbation on each Landau level (as done by Thouless-Kohmoto-Nightingale-den Nijs [TKNdN82]). On the other hand, in the limit of weak fields, one starts with the so-called “tight-binding” Hamiltonian describing the lowest energy sector of a charged particle in a periodic potential, and treats the magnetic field as a perturbation (as done by Harper [Har55] and Hofstadter [Hof76]). These two approaches lead to the same Hamiltonian, with one remarkable exception, as already noted in [TKNdN82]: the value of the magnetic flux is replaced as $\Phi\to 1/\Phi$. As noted in [HKT16], this formal operation coincides precisely with taking the modular dual of the Hamiltonian.
In the case of weak fields, following [Koh89, HKT16] one writes the tight-binding Hamiltonian describing a charged particle in a two-dimensional periodic potential in the usual form $H=T_{x}+T_{x}^{*}+T_{y}+T_{y}^{*}$ where $T_{x}=\sum_{nm}c^{*}_{n+1,m}c_{n,m}$ and $T_{y}=\sum_{nm}c^{*}_{n,m+1}c_{n,m}$ are the translation operators on the lattice along the two axes $x$ and $y$, with $c_{n,m}$ the fermion operators on the lattice at on the position $(x,y)=(n,m)$ with $n,m\in\mathbb{Z}$. We now recall that the translation operator corresponding to a displacement $\vec{r}$ is also $T(\vec{r})=\mathrm{e}^{-\mathrm{i}\vec{r}\cdot\vec{p}}$, which gives $T_{x}=\mathrm{e}^{-\mathrm{i}p_{x}}$, $T_{y}=\mathrm{e}^{-\mathrm{i}p_{y}}$ (we are translating by one lattice site). This gives
@@
The same result can be obtained directly by performing the substitution $\vec{k}\mapsto\vec{p}-\vec{A}$ directly in $E(k_{x},k_{y})$, promoting the energy dispersion to the Hamiltonian operator $H=2\cos{p_{x}}+2\cos{(p_{y}-2\pi\alpha x)}$ ) [Hof76, Koh89]. In the limit of strong fields instead, by treating the periodic potential $V(x,y)$ as a perturbation on each Landau level, one obtains the same Hamiltonian but with the remarkable exception that one now gets $\alpha\to 1/\alpha$, as already mentioned (see [TKNdN82]). Since $[p_{y},x]=[p_{y},p_{x}]=0$, one has that $[p_{y},H]=0$: Hence it is possible to simultaneously diagonalize the Hamiltonian and the momentum $p_{y}$. In particular, for eigenvalues of the momentum $p_{y}\to 0$, the Hamiltonian reduces to the Hamiltonian $H_{\text{Hof}}$ defined in Eq. 8.

### v2

@@
Abstract
The Hofstadter model allows to describe and understand several phenomena in condensed matter such as the quantum Hall effect, Anderson localization, charge pumping, and flat-bands in quasiperiodic structures, and is a rare example of fractality in the quantum world. An apparently unrelated system, the relativistic Toda lattice, has been extensively studied in the context of complex nonlinear dynamics, and more recently for its connection to supersymmetric Yang-Mills theories and topological string theories on Calabi-Yau manifolds in high-energy physics. Here we discuss a recently discovered spectral relationship between the Hofstadter model and the relativistic Toda lattice which has been later conjectured to be related to the Langlands duality of quantum groups. Moreover, by employing similarity transformations compatible with the quantum group structure, we establish a formula parametrizing the energy spectrum of the Hofstadter model in terms of elementary symmetric polynomials and Chebyshev polynomials. The main tools used are the spectral duality of tridiagonal matrices and the representation theory of the elementary quantum group.
quantum Hall effect, Hofstadter model, Aubry-André model, quantum groups, Toda lattice.
@@ Introduction
Among the diversity of natural phenomena, intriguing connections sometimes emerge, where different physical systems share similar mathematical equations or structures. These unexpected parallels may seem coincidental or, conversely, hint at a deeper unity within the physical and mathematical worlds. One such example is the relationship between the relativistic Toda lattice modelToda (1967); Ruijsenaars (1990) in high energy physics and the HofstadterHarper (1955); Hofstadter (1976); Aubry and André (1980) model in condensed matter physics.
@@
The Toda lattice model describes a one-dimensional lattice of $N$ particles with exponentially decaying interactionsToda (1967). Initially conceived as a simple toy model for one-dimensional crystals, it is a prominent example of a nonlinear system in mathematical physics, which stands out for the remarkable property of being exactly solvable and integrable. Moreover, the model exhibits the emergence of solitonic excitations, providing insights into the complex dynamics of nonlinear systems. Furthermore, its relativistic generalizationRuijsenaars (1990) is studied in high-energy physics for its connection to supersymmetric Yang-Mills theories in five dimensionsNekrasov (1998) and topological string theories on Calabi-Yau manifoldsKatz, Klemm, and Vafa (1997); Aganagic et al. (2012); Hatsuda and Mariño (2016); Hatsuda, Katsura, and Tachikawa (2016); Hatsuda, Sugimoto, and Xu (2017).
@@
On the other hand, the Hofstadter model describes noninteracting charged fermions (e.g., electrons) on a two-dimensional lattice in a magnetic field (more generally, a gauge field) perpendicular to the plane or, alternatively, noninteracting fermions in a one-dimensional quasiperiodic latticeHarper (1955); Hofstadter (1976); Aubry and André (1980). The most distinctive feature of the model is the presence of a fractal energy spectrum, which is one of the few examples of fractals in quantum physics, and the very rich phase diagramOsadchy and Avron (2001) with topological gapped phases indexed by a topological invariant, the Chern number of the occupied bandsThouless et al. (1982). The Hofstadter model is the discrete counterpart of a quantum Hall system, in the sense that it describes the Landau levels of free electrons in a magnetic field regularized on a discrete latticeThouless et al. (1982). Indeed, in the weak field limit, i.e., when the magnetic cyclotron radius becomes much larger than the lattice constant, the quantum Hall system and the Hofstadter model become physically equivalent. It is however when the magnetic cyclotron radius is comparable with the lattice constant that the fractal properties emerge, as a consequence of the incommensuration between these two characteristic lengths.
@@
The Hofstadter Hamiltonian is moreover identical to the model introduced by Aubry and AndréAubry and André (1980), describing fermions on a one-dimensional discrete lattice in the presence of a harmonic potential not necessarily commensurate with the lattice parameter. In the Aubry-André model, the role of the magnetic flux per unit cell in the two-dimensional system is played by the angular wavenumber of the harmonic potential (in units of the lattice parameter), and the role of the second dimension is played by the phase shift of the harmonic potential with respect to the discrete lattice, giving rise to a so-called second synthetic dimension. This also leads to the emergence of topological charge pumpingThouless (1983), which shares the same topological origin as the quantum Hall effect. The fractal nature of the spectra emerges here as a consequence of the incommensuration between the periodicity of the harmonic potential and the periodicity of the underlying lattice. This model can describe the effects of disorder, including Anderson localizationAubry and André (1980) and quasiperiodicityKraus et al. (2012); Kraus, Ringel, and Zilberberg (2013); Marra and Nitta (2020).
@@
Physical realizations of the Hofstadter model include microwave photonic systemsKuhl and Stöckmann (1998), Moiré superlattices in graphenePonomarenko et al. (2013); Dean et al. (2013); Hunt et al. (2013), ultracold atoms in incommensurate optical latticesAidelsburger et al. (2013), and interacting phonons in superconducting qubitsRoushan et al. (2017).
@@
Although originating from distinct branches of physics, the $N=2$ relativistic Toda lattice and Hofstadter models are described by Hamiltonians which transform one into the other by a formal substitution of the position and momenta operators $x,p\mapsto\mathrm{i}x,\mathrm{i}p$. Mathematically, this corresponds to substituting Mathieu ($\cos$) and the modified Mathieu ($\cosh$) potentials or operators. Recently, a deep relationship between the energy spectra of the Hofstadter model, the relativistic Toda lattice model, and their modular dual was found by Hatsuda, Katsura, and TachikawaHatsuda, Katsura, and Tachikawa (2016). Building on these results, it has been conjectured that the Hofstadter model is intimately related to the Langlands duality of quantum groupsIkeda (2018), suggesting a deep connection between the Langlands program, quantum geometry, and quasiperiodicity.
@@
We notice that both the Hofstadter and Toda lattice models are known to be related to the root systems $A_{N-1}$. Hence it should be no surprise that $\mathcal{U}_{q}(\mathfrak{sl}_{2})$ (the simplest quantum group) plays a central role in the study of these models. More generally, we stress the fact that noncommutative geometry is the most natural language to investigate the properties of the Hofstadter model and the related Aubry and André model, as well as other quasiperiodic and aperiodic structures emerging in condensed matter physics, such as quasicrystals, Penrose tilings, and disordered systemsBellissard (2003); Putnam (2010); Loring and Hastings (2011); Oyono-Oyono and Petite (2011).
It should be emphasized that the first result below (Theorem II.2) has already been understoodHatsuda, Katsura, and Tachikawa (2016) in the context of the relativistic Toda model, and our contribution is a clarification of the relationship with the characteristic polynomials associated with the Hofstadter model, making use of a “quantum group-adjusted” gauge introduced by Wiegmann and ZabrodinWiegmann and Zabrodin (1994), and further studied by Hatsugai, Kohmoto, and WuHatsugai, Kohmoto, and Wu (1996), and the spectral duality developed by Molinari (see Ref. Molinari, 1997, where general Hamiltonian matrices similar to the one considered here are investigated).
The formula given in Theorem III.9 and the parametrization in Theorem IV.2 are new, and they are obtained by using the representation theory of $\mathcal{U}_{q}(\mathfrak{sl}_{2})$ and standard formulas involving Chebyshev polynomials.
@@
The polynomial relationship of Eq. (1) was first proved by Hatsuda, Katsura, and TachikawaHatsuda, Katsura, and Tachikawa (2016) by comparing the eigenvalue equation of the Hofstadter Hamiltonian $H$, the Toda Hamiltonian $H_{\text{Toda}}$, and its modular dual.
Here, we clarify the argument by using the spectral duality of tridiagonal matrices with corners developed in Ref. Molinari, 1997 and proving that the polynomials involved are indeed the characteristic polynomials determining the spectrum of the Hofstadter model (as observed in Ref. Hatsuda, Katsura, and Tachikawa, 2016, p. 10).
@@
For more details on the result above, see Theorem IV.2 and related discussion above it. This result should be interpreted as follows: in order to gain more insight into the mapping $E\mapsto\tilde{E}$, we can try to obtain explicit formulas for $f$ for all values of $\alpha=P/Q$ and then invert one side of Eq. (1) (at least locally) to get an analytic expression $\tilde{E}=\tilde{E}(E)$. This theorem is an attempt towards this goal. It should be noted that a formula for the Hofstadter spectrum at the mid-band point appeared previously in Ref. Kreft, 1993, by using a strategy related to the one employed in this paper, but without involving the notion of quantum group.
@@
I Preliminaries on mathematical structures and models
@@
There are several such constructions, most notably the Gell-Mann matrices and Sylvester’s shift and clock matricesSylvester (2011). The former are Hermitian and traceless (just like the Pauli matrices), while the latter are unitary and traceless, which makes them preferable for our goals.
Definition I.1.
@@
which justifies the name Weyl-Heisenberg matrices for $V$ and $U$. The Hofstadter model is described by the Hamiltonian $H=V+V^{*}+U+U^{*}$. Before going further, however, we need to introduce some additional mathematical structures. The relation in Eq. (2) defines a universal $C^{*}$-algebra $A_{\omega}$ studied by Rieffel Rieffel (1981), known as noncommutative torus or rotation algebra. It is arguable that the naming “noncommutative torus” should be reserved for the case $\omega=\exp(2\pi\mathrm{i}\alpha)$ and $\alpha$ is irrational, unlike what we assumed so far. This is because when $\alpha$ is rational, $A_{\omega}$ is Morita equivalent to the commutative algebra of continuous functions on the standard torus. By contrast, when $\alpha$ is irrational, the Morita equivalence class is determined by the orbit of the modular action of $\mathrm{PGL}_{2}(\mathbb{Z})$ by linear fractional transformations on $\alpha$. Being these generated by
@@
in the mathematical physics literature $A_{\omega}$ is said to be in modular duality with $A_{\tilde{\omega}}$, where $\tilde{\omega}=\exp(2\pi\mathrm{i}\alpha^{-1})$. Let us remark here that the operator algebras community does not use this naming convention. However, other related dualities, such as Spanier-Whitehead duality in $K$-theoryConnes (1994); Nishikawa and Proietti (2020) or the property of being each other’s von Neumann algebra commutant (with coupling constant $\alpha$)Faddeev (1995), are more frequently mentioned for $A_{\omega}$ and $A_{\tilde{\omega}}$.
Equation 3 provides our second contact point with the Hofstadter model, as those two transformations also appear as the fundamental self-similarity relations of the Hofstadter butterflyHofstadter (1976); Hatsuda, Katsura, and Tachikawa (2016); Ikeda (2018). In addition, the modular duality discussed above has inspired Faddeev’s notion of modular double of a quantum groupFaddeev (2014). The modular double of $\mathcal{U}_{q}(\mathfrak{sl}_{2})$ plays a central role in the study of the relativistic Toda latticeKharchev, Lebedev, and Semenov-Tian-Shansky (2002), particularly in the argument of Ref. Hatsuda, Katsura, and Tachikawa, 2016 where the polynomial relation which inspired this paper is established. To understand this, let us first recall the definition of the elementary quantum group.
Definition I.2.
@@
Following Ref. Kharchev, Lebedev, and Semenov-Tian-Shansky, 2002, we operate a Wick rotation on $\rho_{0},\tilde{\rho}_{0}$, by changing $T_{s}$ into $T_{\mathrm{i}s}$ and $S_{-\mathrm{i}s}$ into $S_{s}$. Note there is a fundamental difference between the representations of the quantum group and those of the rotation algebra: the former are obtained from unbounded representations of the latter. The newly obtained representations will be denoted $\rho,\tilde{\rho}$. Since they still commute, we can combine them with $\phi$ to form a representation of the modular double quantum group
@@
The general definition of this object involves the Langlands dual Lie algebraKharchev, Lebedev, and Semenov-Tian-Shansky (2002). However, we do not need to worry about such details here as $\mathfrak{sl}_{2}$ is self-dual. We can anticipate that the Toda Hamiltonian is $H\propto\rho\phi(E+F)$.
In condensed matter physics, the $N$-particle Toda lattice modelToda (1967) is a seminal toy model describing massive particles on a one-dimensional lattice subject to a nearest-neighbor and exponentially decaying two-body interaction. In the periodic case (particles on a circle) the Hamiltonian reads
@@
with the periodic boundary condition $q_{0}\equiv q_{N}$, $q_{N+1}\equiv q_{1}$, where $p_{n}$ are the canonical momenta and $q_{n}$ the canonical coordinates satisfying the canonical Poisson bracket relations $\{q_{n},p_{m}\}=\delta_{nm}$. This Hamiltonian is nonrelativistic, being invariant up to Galilean transformations but not transformations of the Poincaré group, and is an early example of the family of integrable one-dimensional many-body problems, also called Calogero-Moser systems. A general approach to obtain a relativistic generalization of the Toda lattice model (as well as other Calogero-Moser systems) is to deform the nonrelativistic Hamiltonian to make it compatible with the Poincaré group (see Ref. Ruijsenaars, 1990 for more details). This involves maintaining the nearest-neighbor exponential form of the interaction term, exponentiating the canonical momenta, and combining the resulting terms into an Hamiltonian which will coincide with the time-translation generator of the Poincaré group. Moreover, we want our Hamiltonian to be quantum, that is, to elevate the classical variables to quantum operators satisfying the canonical commutation relation $[q_{n},p_{m}]=\mathrm{i}\hbar\delta_{nm}$. Following this general recipe and using the same approach as Refs. Kharchev, Lebedev, and Semenov-Tian-Shansky, 2002; Hatsuda and Mariño, 2016, we define our $N$-particle relativistic Toda lattice model as
@@
and also gives $[x,p]=\mathrm{i}\hbar/(2\pi\alpha)$. The relationship between quantum mechanics and quantum geometry is thus revealed if one sets $\hbar=2\pi\alpha$, which leads us back to $[x,p]=\mathrm{i}$ and $q=\mathrm{e}^{\mathrm{i}\pi\alpha}$. This Hamiltonian hence is expressed in terms of $\rho(v)$ and $\rho(vu)$, namely the first two generators of the generalized Clifford algebra introduced above (with the difference that these are infinite-dimensional operators). Finally, to uniform our notations to that of Ref. Hatsuda, Katsura, and Tachikawa, 2016, we introduce the canonical transformation $2\pi\alpha x^{\prime}=2\pi\alpha x+p$ (preserving the commutation relation $[x^{\prime},p]=[x,p]=\mathrm{i}$). By doing so and neglecting the prime, this yields the usual form of the relativistic Toda Hamiltonian for $N=2$ particles:
@@
The Hofstadter Hamiltonian describes the Hamiltonian of a charged particle in a magnetic field and in a periodic potential $V(x,y)$ describing a two-dimensional lattice. There are two opposite ways to derive this Hamiltonian: In the limit of strong fields, one takes the Hamiltonian of a charged particle in a field (describing the quantum Hall effect), which is diagonalized in terms of Landau levels, and then treats the periodic potential $V(x,y)$ as a perturbation on each Landau level (as done by Thouless-Kohmoto-Nightingale-den NijsThouless et al. (1982)). On the other hand, in the limit of weak fields, one starts with the so-called “tight-binding” Hamiltonian describing the lowest energy sector of a charged particle in a periodic potential, and treats the magnetic field as a perturbation (as done by HarperHarper (1955) and HofstadterHofstadter (1976)). These two approaches lead to the same Hamiltonian, with one remarkable exception, as already noted in Ref. Thouless et al., 1982: the value of the magnetic flux is replaced as $\Phi\to 1/\Phi$. As noted in Ref. Hatsuda, Katsura, and Tachikawa, 2016, this formal operation coincides precisely with taking the modular dual of the Hamiltonian.
In the case of weak fields, following Ref. Kohmoto, 1989; Hatsuda, Katsura, and Tachikawa, 2016 one writes the tight-binding Hamiltonian describing a charged particle in a two-dimensional periodic potential in the usual form $H=T_{x}+T_{x}^{*}+T_{y}+T_{y}^{*}$ where $T_{x}=\sum_{nm}c^{*}_{n+1,m}c_{n,m}$ and $T_{y}=\sum_{nm}c^{*}_{n,m+1}c_{n,m}$ are the translation operators on the lattice along the two axes $x$ and $y$, with $c_{n,m}$ the fermion operators on the lattice at on the position $(x,y)=(n,m)$ with $n,m\in\mathbb{Z}$. We now recall that the translation operator corresponding to a displacement $\vec{r}$ is also $T(\vec{r})=\mathrm{e}^{-\mathrm{i}\vec{r}\cdot\vec{p}}$, which gives $T_{x}=\mathrm{e}^{-\mathrm{i}p_{x}}$, $T_{y}=\mathrm{e}^{-\mathrm{i}p_{y}}$ (we are translating by one lattice site). This gives
@@
The same result can be obtained directly by performing the substitution $\vec{k}\mapsto\vec{p}-\vec{A}$ directly in $E(k_{x},k_{y})$, promoting the energy dispersion to the Hamiltonian operator $H=2\cos{p_{x}}+2\cos{(p_{y}-2\pi\alpha x)}$ )Hofstadter (1976); Kohmoto (1989). In the limit of strong fields instead, by treating the periodic potential $V(x,y)$ as a perturbation on each Landau level, one obtains the same Hamiltonian but with the remarkable exception that one now gets $\alpha\to 1/\alpha$, as already mentioned (see Ref. Thouless et al., 1982). Since $[p_{y},x]=[p_{y},p_{x}]=0$, one has that $[p_{y},H]=0$: Hence it is possible to simultaneously diagonalize the Hamiltonian and the momentum $p_{y}$. In particular, for eigenvalues of the momentum $p_{y}\to 0$, the Hamiltonian reduces to the Hamiltonian $H_{\text{Hof}}$ defined in Eq. 8.